The rawfonts package

Alan Jeffrey

v0.01

Overview

The rawfonts L^{A}TEX 2ε package is used to provide emulation of L^{A}TEX 2.09 documents which used low-level commands such as \tenrm. If you say:

    \usepackage{rawfonts}

then over sixty fonts such as \tenrm will be loaded into L^{A}TEX. This has a large overhead, for example this document uses:

    8221 words of font info for 30 fonts

but when run with the rawfonts package it uses:

    19294 words of font info for 73 fonts

That is rawfonts can cause your document to load twice as many fonts. (This overhead is why L^{A}TEX 2ε does not define \tenrm and friends by default.)

If you want to only load a small number of fonts, you can use the only option, for example to only load \tenrm and \tensf:

    \usepackage[only,tenrm,tensf]{rawfonts}

The rawfonts package is intended mainly for use with L^{A}TEX 2.09 documents, and you might want to have the package loaded every time you use L^{A}TEX 2ε in compatibility mode. In this case, you should add the line:

    \RequirePackage{rawfonts}

to your latex209.cfg L^{A}TEX 2.09 compatibility configuration file.

Implementation

The driver for the documentation you’re now reading.

1 ⟨∗driver⟩

 2 \documentclass{ltxdoc}
 3 \begin{document}
 4 \DocInput{rawfonts.dtx}
 5 \end{document}

6 ⟨/driver⟩

This is a L^{A}TEX 2ε package.

7 ⟨∗package⟩

 8 \NeedsTeXFormat{LaTeX2e}
 9 \ProvidesPackage{rawfonts}
10 [1994/05/08 Low-level LaTeX 2.09 font compatibility]

The rawfonts package uses the somedefs package.

11 \RequirePackage{somedefs}

By default, all the fonts are loaded, but the only option says that only the ones specified as options should be.

12 \UseAllDefinitions
13 \DeclareOption{only}{\UseSomeDefinitions}
14 \DeclareOption*{\UseDefinition{\CurrentOption}}
15 \ProcessOptions

The rest of the code loads fonts. Five point:

16 \ProvidesDefinition{\DeclareFixedFont{\fivrm}{OT1}{cmr}{m}{n}{\@vpt}}
17 \ProvidesDefinition{\DeclareFixedFont{\fivmi}{OML}{cmm}{m}{it}{\@vpt}}
18 \ProvidesDefinition{\DeclareFixedFont{\fivsy}{OMS}{cmsy}{m}{n}{\@vpt}}
19 \ProvidesDefinition{\DeclareFixedFont{\fivly}{U}{lasy}{m}{n}{\@vpt}}

Six point:

20 \ProvidesDefinition{\DeclareFixedFont{\sixrm}{OT1}{cmr}{m}{n}{\@vipt}}
21 \ProvidesDefinition{\DeclareFixedFont{\sixmi}{OML}{cmm}{m}{it}{\@vipt}}
22 \ProvidesDefinition{\DeclareFixedFont{\sixsy}{OMS}{cmsy}{m}{n}{\@vipt}}
23 \ProvidesDefinition{\DeclareFixedFont{\sixly}{U}{lasy}{m}{n}{\@vipt}}

Seven point:

24 \ProvidesDefinition{\DeclareFixedFont{\sevrm}{OT1}{cmr}{m}{n}{\@viipt}}
25 \ProvidesDefinition{\DeclareFixedFont{\sevmi}{OML}{cmm}{m}{it}{\@viipt}}
26 \ProvidesDefinition{\DeclareFixedFont{\sevsy}{OMS}{cmsy}{m}{n}{\@viipt}}
27 \ProvidesDefinition{\DeclareFixedFont{\sevit}{OT1}{cmr}{m}{it}{\@viipt}}
28 \ProvidesDefinition{\DeclareFixedFont{\sevly}{U}{lasy}{m}{n}{\@viipt}}

Eight point:

29 \ProvidesDefinition{\DeclareFixedFont{\egtrm}{OT1}{cmr}{m}{n}{\@viiipt}}
30 \ProvidesDefinition{%
31  \DeclareFixedFont{\egtmi}{OML}{cmm}{m}{it}{\@viiipt}}
32 \ProvidesDefinition{%
33  \DeclareFixedFont{\egtsy}{OMS}{cmsy}{m}{n}{\@viiipt}}
34 \ProvidesDefinition{%

35  \DeclareFixedFont{\egtit}{OT1}{cmr}{m}{it}{\@viiipt}}
36 \ProvidesDefinition{\DeclareFixedFont{\egtly}{U}{lasy}{m}{n}{\@viiipt}}

Nine point:

37 \ProvidesDefinition{\DeclareFixedFont{\ninrm}{OT1}{cmr}{m}{n}{\@ixpt}}
38 \ProvidesDefinition{\DeclareFixedFont{\ninmi}{OML}{cmm}{m}{it}{\@ixpt}}
39 \ProvidesDefinition{\DeclareFixedFont{\ninsy}{OMS}{cmsy}{m}{n}{\@ixpt}}
40 \ProvidesDefinition{\DeclareFixedFont{\ninit}{OT1}{cmr}{m}{it}{\@ixpt}}
41 \ProvidesDefinition{\DeclareFixedFont{\ninbf}{OT1}{cmr}{bx}{n}{\@ixpt}}
42 \ProvidesDefinition{\DeclareFixedFont{\nintt}{OT1}{cmtt}{m}{n}{\@ixpt}}
43 \ProvidesDefinition{\DeclareFixedFont{\ninly}{U}{lasy}{m}{n}{\@ixpt}}

Ten point:

44 \ProvidesDefinition{\DeclareFixedFont{\tenrm}{OT1}{cmr}{m}{n}{\@xpt}}
45 \ProvidesDefinition{\DeclareFixedFont{\tenmi}{OML}{cmm}{m}{it}{\@xpt}}
46 \ProvidesDefinition{\DeclareFixedFont{\tensy}{OMS}{cmsy}{m}{n}{\@xpt}}
47 \ProvidesDefinition{\DeclareFixedFont{\tenit}{OT1}{cmr}{m}{it}{\@xpt}}
48 \ProvidesDefinition{\DeclareFixedFont{\tensl}{OT1}{cmr}{m}{sl}{\@xpt}}
49 \ProvidesDefinition{\DeclareFixedFont{\tenbf}{OT1}{cmr}{bx}{n}{\@xpt}}
50 \ProvidesDefinition{\DeclareFixedFont{\tentt}{OT1}{cmtt}{m}{n}{\@xpt}}
51 \ProvidesDefinition{\DeclareFixedFont{\tensf}{OT1}{cmss}{m}{n}{\@xpt}}
52 \ProvidesDefinition{\DeclareFixedFont{\tenly}{U}{lasy}{m}{n}{\@xpt}}
53 \ProvidesDefinition{\DeclareFixedFont{\tenex}{OMX}{cmex}{m}{n}{\@xpt}}

Eleven point:

54 \ProvidesDefinition{\DeclareFixedFont{\elvrm}{OT1}{cmr}{m}{n}{\@xipt}}
55 \ProvidesDefinition{\DeclareFixedFont{\elvmi}{OML}{cmm}{m}{it}{\@xipt}}
56 \ProvidesDefinition{\DeclareFixedFont{\elvsy}{OMS}{cmsy}{m}{n}{\@xipt}}
57 \ProvidesDefinition{\DeclareFixedFont{\elvit}{OT1}{cmr}{m}{it}{\@xipt}}
58 \ProvidesDefinition{\DeclareFixedFont{\elvsl}{OT1}{cmr}{m}{sl}{\@xipt}}
59 \ProvidesDefinition{\DeclareFixedFont{\elvbf}{OT1}{cmr}{bx}{n}{\@xipt}}
60 \ProvidesDefinition{\DeclareFixedFont{\elvtt}{OT1}{cmtt}{m}{n}{\@xipt}}
61 \ProvidesDefinition{\DeclareFixedFont{\elvsf}{OT1}{cmss}{m}{n}{\@xipt}}
62 \ProvidesDefinition{\DeclareFixedFont{\elvly}{U}{lasy}{m}{n}{\@xipt}}

Twelve point:

63 \ProvidesDefinition{\DeclareFixedFont{\twlrm}{OT1}{cmr}{m}{n}{\@xiipt}}
64 \ProvidesDefinition{\DeclareFixedFont{\twlmi}{OML}{cmm}{m}{it}{\@xiipt}}
65 \ProvidesDefinition{\DeclareFixedFont{\twlsy}{OMS}{cmsy}{m}{n}{\@xiipt}}
66 \ProvidesDefinition{\DeclareFixedFont{\twlit}{OT1}{cmr}{m}{it}{\@xiipt}}
67 \ProvidesDefinition{\DeclareFixedFont{\twlsl}{OT1}{cmr}{m}{sl}{\@xiipt}}
68 \ProvidesDefinition{\DeclareFixedFont{\twlbf}{OT1}{cmr}{bx}{n}{\@xiipt}}
69 \ProvidesDefinition{\DeclareFixedFont{\twltt}{OT1}{cmtt}{m}{n}{\@xiipt}}
70 \ProvidesDefinition{\DeclareFixedFont{\twlsf}{OT1}{cmss}{m}{n}{\@xiipt}}
71 \ProvidesDefinition{\DeclareFixedFont{\twlly}{U}{lasy}{m}{n}{\@xiipt}}

Fourteen point:

72 \ProvidesDefinition{\DeclareFixedFont{\frtnrm}{OT1}{cmr}{m}{n}{\@xivpt}}
73 \ProvidesDefinition{%
74  \DeclareFixedFont{\frtnmi}{OML}{cmm}{m}{it}{\@xivpt}}
75 \ProvidesDefinition{%
76  \DeclareFixedFont{\frtnsy}{OMS}{cmsy}{m}{n}{\@xivpt}}

77 \ProvidesDefinition{%
78  \DeclareFixedFont{\frtnbf}{OT1}{cmr}{bx}{n}{\@xivpt}}
79 \ProvidesDefinition{\DeclareFixedFont{\frtnly}{U}{lasy}{m}{n}{\@xivpt}}

Seventeen point:

80 \ProvidesDefinition{%
81  \DeclareFixedFont{\svtnrm}{OT1}{cmr}{m}{n}{\@xviipt}}
82 \ProvidesDefinition{%
83  \DeclareFixedFont{\svtnmi}{OML}{cmm}{m}{it}{\@xviipt}}
84 \ProvidesDefinition{%
85  \DeclareFixedFont{\svtnsy}{OMS}{cmsy}{m}{n}{\@xviipt}}
86 \ProvidesDefinition{%
87  \DeclareFixedFont{\svtnbf}{OT1}{cmr}{bx}{n}{\@xviipt}}
88 \ProvidesDefinition{\DeclareFixedFont{\svtnly}{U}{lasy}{m}{n}{\@xviipt}}

Twenty point:

89 \ProvidesDefinition{\DeclareFixedFont{\twtyrm}{OT1}{cmr}{m}{n}{\@xxpt}}
90 \ProvidesDefinition{\DeclareFixedFont{\twtymi}{OML}{cmm}{m}{it}{\@xxpt}}
91 \ProvidesDefinition{\DeclareFixedFont{\twtysy}{OMS}{cmsy}{m}{n}{\@xxpt}}
92 \ProvidesDefinition{\DeclareFixedFont{\twtyly}{U}{lasy}{m}{n}{\@xxpt}}

Twenty-five point:

93 \ProvidesDefinition{\DeclareFixedFont{\twfvrm}{OT1}{cmr}{m}{n}{\@xxvpt}}

That’s it.

94 ⟨/package⟩

Index

Numbers written in italic refer to the page where the corresponding entry is described, the ones underlined to the code line of the definition, the rest to the code lines where the entry is used.

Symbols

\@ixpt . . . . . . . . . 37–43 \@viiipt 29, 31, 33, 35, 36 \@viipt . . . . . . . . 24–28 \@vipt . . . . . . . . . 20–23 \@vpt . . . . . . . . . . 16–19 \@xiipt . . . . . . . . 63–71 \@xipt . . . . . . . . . 54–62 \@xivpt 72, 74, 76, 78, 79 \@xpt . . . . . . . . . . 44–53 \@xviipt 81, 83, 85, 87, 88 \@xxpt . . . . . . . . . 89–92 \@xxvpt . . . . . . . . . . 93

B

\begin . . . . . . . . . . . . 3

C

                          \CurrentOption . . . . 14

D

                          \DeclareFixedFont .

. . . . . . 16–29, 31, 33, 35–72, 74, 76, 78, 79, 81, 83, 85, 87–93

                          \DeclareOption . . 13, 14

\DocInput . . . . . . . . . 4

                          \documentclass . . . . . 2

E \egtit . . . . . . . . . . . 35 \egtly . . . . . . . . . . . 36 \egtmi . . . . . . . . . . . 31 \egtrm . . . . . . . . . . . 29 \egtsy . . . . . . . . . . . 33 \elvbf . . . . . . . . . . . 59 \elvit . . . . . . . . . . . 57 \elvly . . . . . . . . . . . 62 \elvmi . . . . . . . . . . . 55 \elvrm . . . . . . . . . . . 54 \elvsf . . . . . . . . . . . 61 \elvsl . . . . . . . . . . . 58 \elvsy . . . . . . . . . . . 56 \elvtt . . . . . . . . . . . 60 \end . . . . . . . . . . . . . . 5

F \fivly . . . . . . . . . . . 19 \fivmi . . . . . . . . . . . 17 \fivrm . . . . . . . . . . . 16 \fivsy . . . . . . . . . . . 18 \frtnbf . . . . . . . . . . 78 \frtnly . . . . . . . . . . 79 \frtnmi . . . . . . . . . . 74 \frtnrm . . . . . . . . . . 72 \frtnsy . . . . . . . . . . 76

N

\NeedsTeXFormat . . . . 8

\ninbf . . . . . . . . . . . 41 \ninit . . . . . . . . . . . 40 \ninly . . . . . . . . . . . 43 \ninmi . . . . . . . . . . . 38 \ninrm . . . . . . . . . . . 37 \ninsy . . . . . . . . . . . 39 \nintt . . . . . . . . . . . 42

P

\ProcessOptions . . . 15
\ProvidesDefinition

. . . . . . 16–30, 32, 34, 36–73, 75, 77, 79, 80, 82, 84, 86, 88–93

\ProvidesPackage . . . 9

R

\RequirePackage . . . 11

S

\sevit . . . . . . . . . . . 27 \sevly . . . . . . . . . . . 28 \sevmi . . . . . . . . . . . 25 \sevrm . . . . . . . . . . . 24 \sevsy . . . . . . . . . . . 26 \sixly . . . . . . . . . . . 23 \sixmi . . . . . . . . . . . 21 \sixrm . . . . . . . . . . . 20 \sixsy . . . . . . . . . . . 22 \svtnbf . . . . . . . . . . 87 \svtnly . . . . . . . . . . 88 \svtnmi . . . . . . . . . . 83 \svtnrm . . . . . . . . . . 81 \svtnsy . . . . . . . . . . 85

T

\tenbf . . . . . . . . . . . 49 \tenex . . . . . . . . . . . 53 \tenit . . . . . . . . . . . 47 \tenly . . . . . . . . . . . 52 \tenmi . . . . . . . . . . . 45 \tenrm . . . . . . . . . . . 44 \tensf . . . . . . . . . . . 51 \tensl . . . . . . . . . . . 48 \tensy . . . . . . . . . . . 46 \tentt . . . . . . . . . . . 50 \twfvrm . . . . . . . . . . 93 \twlbf . . . . . . . . . . . 68 \twlit . . . . . . . . . . . 66 \twlly . . . . . . . . . . . 71 \twlmi . . . . . . . . . . . 64 \twlrm . . . . . . . . . . . 63 \twlsf . . . . . . . . . . . 70 \twlsl . . . . . . . . . . . 67 \twlsy . . . . . . . . . . . 65 \twltt . . . . . . . . . . . 69 \twtyly . . . . . . . . . . 92 \twtymi . . . . . . . . . . 90 \twtyrm . . . . . . . . . . 89 \twtysy . . . . . . . . . . 91

U

\UseAllDefinitions . 12
\UseDefinition . . . . 14
\UseSomeDefinitions 13